# Legilimens: Practical and Unified Content Moderation for Large Language Model Services

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 31th ACM Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, Utahisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning
@@ 4.1.1. Conceptual Feature Extraction
(8) ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{X})\triangleq\left[\widetilde{\mathcal{H}}_{(K-m)}(\mathbf{X})\oplus\cdots\oplus\widetilde{\mathcal{H}}_{K}(\mathbf{X})\right].$
The reason for deriving conceptual features from the last several blocks is to utilize more capacity of the host LLM, given that ~$\widetilde{\mathcal{H}}_{(K-m):K}$ can be viewed as containing all parameters of the host LLM. Note that, no additional computation is required to obtain the side features in Equation (8).
@@ 4.1.2. Probing Position
As shown in Figure 2, the attention span during the decoding of the first token $r_{1}$ encompasses the entire prompt, indicating that the conceptual features derived in this process, ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P})$, encapsulate a summarized knowledge of prompt $\mathbf{p}$. Similarly, the attention span during the decoding the last token [eos] encompasses both the prompt and the response, encapsulating a summarized knowledge of prompt $\mathbf{p}$ and response $\mathbf{r}$ in ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P}\oplus\mathbf{R})$. Note that, both ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P})$ and ~$\widetilde{\mathcal{H}}_{(K-m):K}(\mathbf{P}\oplus\mathbf{R})$ are of a constant size of $m\times d_{\text{model}}$, where $d_{\text{model}}$ is the output dimension of Transformer blocks. Reasoning on these two conceptual features only introduces constant overhead.
@@ 4.2.1. Architecture
In this part, we construct a lightweight classifier to materialize $\mathcal{C}_{\theta\backslash\psi}$. Based on the comprehensive representation summarized by the host LLM, we find that a simple multi-layer perceptron (MLP) with few parameters is adequate for accurate content moderation, denoted as $\mathcal{C}_{\theta\backslash\psi}:\mathbbm{R}^{m\times d_{\text{model}}}\rightarrow\mathbbm{Y}$. Therefore, the entire moderator is composed in this way: $\mathcal{M}_{\theta}=\mathcal{C}_{\theta\backslash\psi}\circ\widetilde{\mathcal{H}}_{(K-m):K}$. We validate the effectiveness of our lightweight moderator in §5.2 compared with five baselines.
@@ Acknowledgement
We sincerely thank the anonymous reviewers for their valuable comments and dedication. This work was supported by Ant Group through CCF-Ant Research Fund. Yanjiao Chen is the corresponding author.

### v2

@@ Abstract.
Content moderation; large language model; jailbreaking
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3658644.3690322conference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAisbn: 979-8-4007-0636-3/24/10ccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Artificial intelligence
@@ 4.1.1. Conceptual Feature Extraction
(8) ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{X})\triangleq\left[\widetilde{\mathcal{H}}_{(K-m+1)}(\mathbf{X})\oplus\cdots\oplus\widetilde{\mathcal{H}}_{K}(\mathbf{X})\right].$
The reason for deriving conceptual features from the last several blocks is to utilize more capacity of the host LLM, given that ~$\widetilde{\mathcal{H}}_{[-m:]}$ can be viewed as containing all parameters of the host LLM. Note that, no additional computation is required to obtain the side features in Equation (8).
@@ 4.1.2. Probing Position
As shown in Figure 2, the attention span during the decoding of the first token $r_{1}$ encompasses the entire prompt, indicating that the conceptual features derived in this process, ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{P})$, encapsulate a summarized knowledge of prompt $\mathbf{p}$. Similarly, the attention span during the decoding the last token [eos] encompasses both the prompt and the response, encapsulating a summarized knowledge of prompt $\mathbf{p}$ and response $\mathbf{r}$ in ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{P}\oplus\mathbf{R})$. Note that, both ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{P})$ and ~$\widetilde{\mathcal{H}}_{[-m:]}(\mathbf{P}\oplus\mathbf{R})$ are of a constant size of $m\times d_{\text{model}}$, where $d_{\text{model}}$ is the output dimension of Transformer blocks. Reasoning on these two conceptual features only introduces constant overhead.
@@ 4.2.1. Architecture
In this part, we construct a lightweight classifier to materialize $\mathcal{C}_{\theta\backslash\psi}$. Based on the comprehensive representation summarized by the host LLM, we find that a simple multi-layer perceptron (MLP) with few parameters is adequate for accurate content moderation, denoted as $\mathcal{C}_{\theta\backslash\psi}:\mathbbm{R}^{m\times d_{\text{model}}}\rightarrow\mathbbm{Y}$. Therefore, the entire moderator is composed in this way: $\mathcal{M}_{\theta}=\mathcal{C}_{\theta\backslash\psi}\circ\widetilde{\mathcal{H}}_{[-m:]}$. We validate the effectiveness of our lightweight moderator in §5.2 compared with five baselines.
@@ Acknowledgement
We sincerely thank the anonymous reviewers for their valuable comments and dedication. This work was supported by China NSFC Grant 61925109 and by Ant Group through CCF-Ant Research Fund. Yanjiao Chen is the corresponding author.